# On regions of mixed unitarity for semigroups of unital quantum channels

### v1

@@ 3. Discrete semigroups of mixed unitary channels
Note that $\mathrm{Ad}_{U}$ is an automorphism of $\mathbb{M}_{d}$ and it leaves $\mathcal{P}(\Phi)$ and $\mathcal{N}(\Phi)$ invariant. Taking $\gamma=\mathrm{Ad}_{U^{*}}$, we see that $\gamma\circ\Phi=\mathbf{E}_{\mathcal{P}}+\mathrm{Ad}_{U^{*}}\circ\beta$. Note that $\mathrm{Ad}_{U^{*}}$ being an automorphism,
$\|(\mathrm{Ad}_{U^{*}}\circ\beta)^{n}\|=\|\beta^{n}\|$
converges to zero. Consequently $(\gamma\circ\Phi)^{n}$ converges to $\mathbf{E}_{\mathcal{P}}$. ∎
@@
We consider the decomposition $\Phi=\alpha+\beta$ and other notation as in Theorem 3.8. Note that $\alpha^{n}=\Phi^{n}\circ\mathbf{E}_{\mathcal{P}}$ and $\mathrm{Ad}_{(U^{*})^{n}}\circ\alpha^{n}=\mathbf{E}_{\mathcal{P}}$ for all $n\geq 1$. Thus by taking $\tau_{n}:=\mathrm{Ad}_{(U^{*})^{n}}\circ\Phi^{n}$ and $X\in\mathcal{P}(\Phi)$, we get
@@
That is, $\tau_{n}\in\mathcal{F}(\mathcal{P}(\Phi))$ for all $n\geq 1$. Furthermore,
@@
This implies that $\tau_{n}\to\mathbf{E}_{\mathcal{P}}$ as $n\to\infty$. Consider $r_{\mathcal{P}(\Phi)}>0$ as in Theorem 3.4, then we get a $N\geq 1$ such that $\left\lVert\tau^{n}-\mathbf{E}_{\mathcal{P}}\right\rVert\leq r_{\mathcal{P}(\Phi)}$ for all $n\geq N$. Now, by Theorem 3.4, $\tau_{n}$ is mixed unitary for all $n\geq N$. Consequently, $\Phi^{n}=\mathrm{Ad}_{U^{n}}\circ\tau_{n}$ is mixed unitary for all $n\geq N$. ∎
@@ 4. Quantum dynamical semigroups of unital quantum channels
Theorem 3.9 establishes that any discrete semigroup of unital quantum channels is eventually mixed unitary. A natural extension is to ask whether an analogous result holds for continuous semigroups of unital quantum channels. To this end, we start by recalling the relevant definitions and results for quantum dynamical semigroups.
@@
is a well-defined (bounded) linear map on $\mathbb{M}_{d}$, called the (infinitesimal generator of the semigroup $\{\Phi_{t}\}_{t\geq 0}$. Furthermore,
@@
The generators of quantum dynamical semigroups is characterized by the famous GKLS theorem (See [GKS76, Lin76, KRP, CP17]), namely, we have the following theorem:
@@
[GKS76] Consider a one parameter semigroup of linear maps $\{\Phi_{t}\}_{t\geq 0}$ on $\mathbb{M}_{d}$. Then the following are equivalent:
@@
In this article we are only considering unital quantum channels. It follows from the previous theorem that a linear map $\mathcal{L}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is the generator of a one parameter semigroup of unital quantum channels iff it has the form:
@@
In the case of the continuous dynamical semigroup of mixed unitary channels, the characterization of generators is known in [KuMa87]. They also gave an example of a (continuous) dynamical semigroup of unital quantum channels that is not mixed unitary for some time (see Example 5.10). So, not every dynamical semigroup of unital quantum channels is mixed unitary for all times. Naturally, one can ask what is the behaviour of this semigroup at infinity. More precisely, we are interested in the question: For a given (continuous) dynamical semigroup $\{\Phi_{t}\}_{t\geq 0}$ of unital quantum channels on $\mathbb{M}_{d}$, does there exists a time $t_{0}>0$ such that for all times $t\geq t_{0}$, $\Phi_{t}$ becomes mixed unitary? In this article, we prove that the answer to the above question is affirmative (see Theorem 4.7).
@@
[KuMa87] Let $\{\Phi_{t}\}_{t\geq 0}$ be a quantum dynamical semigroup of unital channels. Then the following are equivalent:
@@
    where $H,L_{j}$ are Hermitian, $U_{k}$ are unitary, $\lambda_{k}\geq 0$ with $\sum_{k=1}^{\ell}\lambda_{k}=1$ and $n,\ell\geq 0$;
@@
In the following, for given $\lambda\in\mathbb{C}$, the Jordan block of size $r\in\mathbb{N}$ with eigenvalue $\lambda$ is defined and denoted as
$J_{r}(\lambda)=\begin{bmatrix}\lambda&1&0&\cdots&0&0\\ 0&\lambda&1&\cdots&0&0\\ \vdots&\vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&0&\cdots&\lambda&1\\ 0&0&0&\cdots&0&\lambda\end{bmatrix}\in\mathbb{M}_{r}.$
Note that for $r=1$, $J_{1}(\lambda)=[\lambda]\in\mathbb{M}_{1}$. The following Lemma is perhaps known in the literature, but we are providing a proof for completeness.
@@
Let $\mathcal{L}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be a linear map and let $\bigoplus_{k=1}^{m}J_{d_{k}}(\lambda_{k})$ be the Jordan decomposition of $\mathcal{L}$ w.r.t. some basis of $\mathbb{M}_{d}$. If $e^{t\mathcal{L}}$ is a quantum dynamical semigroup of unital maps on $\mathbb{M}_{d}$,
@@
$e^{tJ_{d_{k}}(\lambda)}$ $=e^{t\lambda I_{d_{k}}+N_{k}}$
$=e^{t\lambda I_{d_{k}}}e^{tN_{k}}$
$=e^{t\lambda}I_{d_{k}}\big(I_{d_{k}}+tN_{k}+\frac{t^{2}N_{k}^{2}}{2!}+\cdots+\frac{t^{d_{k}-1}N_{k}^{d_{k}-1}}{(d_{k}-1)!}\big).$
@@
Let $\mathcal{L}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be the generator of a quantum dynamical semigroup of unital channels. Then the space $\mathcal{P}:=\operatorname{\textsf{span}}\{X\in\mathbb{M}_{d}:\mathcal{L}(X)=iaX,\quad a\in\mathbb{R}\}$ is a $C^{*}$-sub algebra of $\mathbb{M}_{d}$.
@@
To prove $\mathcal{P}$ is a $C^{*}$-subalgebra of $\mathbb{M}_{d}$, it is enough to show that it is closed under multiplication and taking adjoints. Let $X,Y\in\mathbb{M}_{d}$ and $a,b\in\mathbb{R}$ be such that $\mathcal{L}(X)=iaX$ and $\mathcal{L}(Y)=ibY$. Then for all $t\geq 0$,
@@
This shows that $X,Y$ is in the peripheral space of $e^{t\mathcal{L}}$. Now, as the peripheral space of $e^{t\mathcal{L}}$ lies in the multiplicative domain of $e^{t\mathcal{L}}$, we get that
@@
By differentiating both sides at $t=0$, we obtain $\mathcal{L}(XY)=i(a+b)XY$. This shows that it is closed under multiplication. Now, $\mathcal{P}$ is closed under taking adjoints because $\mathcal{L}$ is *-preserving. This completes the proof. ∎
The following is a one parameter semigroup version of Theorem 3.1 of [BhSa23]. For more on the peripheral Poisson boundary of quantum dynamical semigroups one may see [BD25].
@@
    $\mathbb{M}_{d}=\mathcal{P}\oplus\mathcal{N}$, where $\mathcal{N}=\{Y:\lim_{t\to\infty}e^{t\mathcal{L}}(Y)=0\}$;
@@
    $\Phi_{t}=\alpha_{t}+\beta_{t}$, where $\alpha_{t}:=\Phi_{t}\circ\mathbf{E}_{\mathcal{P}}:\mathbb{M}_{d}\to\mathcal{P}\subseteq\mathbb{M}_{d}$ where $\mathbf{E}_{\mathcal{P}}$ is the unique trace-preserving conditional expectation onto $\mathcal{P}$ and $\beta_{t}:=\Phi_{t}-\alpha_{t}:\mathbb{M}_{d}\to\mathcal{N}\subseteq\mathbb{M}_{d}$ satisfies $\lim_{t\to\infty}\beta_{t}=0$.
@@
$(i)$ Using the Jordan decomposition of the generator $\mathcal{L}$ of $\{\Phi_{t}\}$, we write
@@
where $\sigma(\mathcal{L})=\{\lambda_{1},\lambda_{2},\cdots,\lambda_{n}\}$ is the set of eigenvalues of $\mathcal{L}$ and $\nu(\lambda_{i})$ is the algebraic multiplicity of the eigenvalue $\lambda_{i}$. From Lemma 4.4, if the real part of $\lambda_{i}=0$, all the Jordan blocks corresponding to the eigenvalue $\lambda_{i}$ are of dimension one and hence the algebraic multiplicity is the same as its geometric multiplicity. Therefore,
@@
where $\sigma_{pi}(\mathcal{L})$ be the set of all purely imaginary eigenvalues of $\mathcal{L}$. Now, if $\lambda_{k}\in\sigma(L)$ be such that real part of $\lambda_{k}\neq 0$, then again by Lemma 4.4, we have $e^{tJ_{k}(\lambda_{k})}\to 0$ as $t\to\infty$. Thus, for $Y\in\textsf{ker}(\mathcal{L}-\lambda_{i}\operatorname{\textsf{id}})^{\nu(\lambda_{i})}$, we have $e^{t\mathcal{L}}(Y)\to 0$ as $t\to\infty$. By letting $\mathcal{N}:=\bigoplus_{\lambda\in\sigma(\mathcal{L})\setminus\sigma_{pi}(\mathcal{L})}\textsf{ker}(\mathcal{L}-\lambda\operatorname{\textsf{id}})^{\nu(\lambda)}$ we get the required decomposition.
@@
$\operatorname{\textsf{tr}}(X^{*}X)=\operatorname{\textsf{tr}}(e^{t\mathcal{L}}(X^{*}X))=\operatorname{\textsf{tr}}(e^{t\mathcal{L}}(X^{*})e^{t\mathcal{L}}(X))=0.$

### v2

@@ 3. Discrete semigroups of mixed unitary channels
Note that $\mathrm{Ad}_{U}$ is an automorphism of $\mathbb{M}_{d}$ and it leaves $\mathcal{P}$ and $\mathcal{N}$ invariant. ∎
@@
We consider the decomposition $\Phi=\alpha+\beta$ and other notation as in Theorem 3.8. Note that $\alpha^{n}=\Phi^{n}\circ\mathbf{E}_{\mathcal{P}}$ and $\mathrm{Ad}_{(U^{*})^{n}}\circ\alpha^{n}=\mathbf{E}_{\mathcal{P}}$ for all $n\geq 1$. Thus by taking $\tau_{n}:=\mathrm{Ad}_{(U^{*})^{n}}\circ\Phi^{n}$ and $X\in\mathcal{P}$, we get
@@
That is, $\tau_{n}\in\mathcal{F}(\mathcal{P})$ for all $n\geq 1$. Furthermore,
@@
This implies that $\tau_{n}\to\mathbf{E}_{\mathcal{P}}$ as $n\to\infty$. Consider $r_{\mathcal{P}}>0$ as in Theorem 3.4, then we get a $N\geq 1$ such that $\left\lVert\tau_{n}-\mathbf{E}_{\mathcal{P}}\right\rVert\leq r_{\mathcal{P}}$ for all $n\geq N$. Now, by Theorem 3.4, $\tau_{n}$ is mixed unitary for all $n\geq N$. Consequently, $\Phi^{n}=\mathrm{Ad}_{U^{n}}\circ\tau_{n}$ is mixed unitary for all $n\geq N$. ∎
@@ 4. Quantum dynamical semigroups of unital quantum channels
Theorem 3.9 establishes that every discrete semigroup of unital quantum channels is eventually mixed unitary. Naturally, this raises the question of whether an analogous result holds for continuous semigroups. To this end, we start by recalling the relevant definitions and results for quantum dynamical semigroups.
@@
is a well-defined linear map on $\mathbb{M}_{d}$, called the infinitesimal generator of the semigroup $\{\Phi_{t}\}_{t\geq 0}$. Furthermore,
@@
The generators of quantum dynamical semigroups is characterized by the famous GKLS theorem (See [15, 28, 31, 12]), namely, we have the following theorem:
@@
[15] Consider a one parameter semigroup of linear maps $\{\Phi_{t}\}_{t\geq 0}$ on $\mathbb{M}_{d}$. Then the following are equivalent:
@@
In this article, we restrict our focus to unital quantum channels. In view of the preceding theorem, it follows that a linear map $\mathcal{L}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ is the generator of a quantum dynamical semigroup of unital quantum channels iff it has the form:
@@
The operators $H,L_{j}$’s appearing in equation (4.1) are called as GKLS coefficients/parameters of $\mathcal{L}$. They are not uniquely determined. However, the extent of non-uniqueness can be described (See page 272 of [31]). We simply choose and fix one representation of $\mathcal{L}$ as in (4.1) and work with it.
The generators of continuous dynamical semigroups of mixed unitary channels are characterized in [25], where it is also shown that in general unital semigroups of quantum channels are not necessarily mixed unitary at all times (See Example 5.10). This raises the question of their asymptotic behavior: for a given unital dynamical semigroup $\{\Phi_{t}\}_{t\geq 0}$ on $\mathbb{M}_{d}$, must there exist a threshold $t_{0}$ after which $\Phi_{t}$ remains mixed unitary? In Theorem 4.12, we prove that this is indeed the case.
@@
[25] Let $\{\Phi_{t}\}_{t\geq 0}$ be a quantum dynamical semigroup of unital channels. Then the following are equivalent:
@@
    where $H,L_{j}$ are Hermitian, $U_{k}$ are unitary, $\lambda_{k}\geq 0$ and $n,\ell\geq 0$;
@@
We analyze the semigroup using the Jordan decomposition of its generator. For a given $\lambda\in\mathbb{C}$, let $J_{r}(\lambda)\in\mathbb{M}_{r}$ denote the Jordan block of size $r\in\mathbb{N}$ associated with the eigenvalue $\lambda$. This matrix is defined by having $\lambda$ at each diagonal entry and $1$ at each entry of the super-diagonal. Note that for $r=1$, $J_{1}(\lambda)=[\lambda]\in\mathbb{M}_{1}$. The following Lemma is perhaps known in the literature, but we are providing a proof for completeness.
@@
Let $\mathcal{L}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be a linear map and let $\bigoplus_{k=1}^{m}J_{d_{k}}(\lambda_{k})$ be the Jordan decomposition of $\mathcal{L}$ with respect to some basis of $\mathbb{M}_{d}$. If $e^{t\mathcal{L}}$ is a quantum dynamical semigroup of unital maps on $\mathbb{M}_{d}$,
@@
$e^{tJ_{d_{k}}(\lambda)}$ $=e^{t\lambda}I_{d_{k}}\big(I_{d_{k}}+tN_{k}+\frac{t^{2}N_{k}^{2}}{2!}+\cdots+\frac{t^{d_{k}-1}N_{k}^{d_{k}-1}}{(d_{k}-1)!}\big).$
@@
Let $\mathcal{L}:\mathbb{M}_{d}\to\mathbb{M}_{d}$ be the generator of a quantum dynamical semigroup of unital channels. Then the space $\mathcal{P}:=\operatorname{\textsf{span}}\{X\in\mathbb{M}_{d}:\mathcal{L}(X)=iaX,\quad a\in\mathbb{R}\}$ is a unital $C^{*}$-sub algebra of $\mathbb{M}_{d}$.
@@
Clearly $\mathcal{P}$ contains the identity. As $\mathcal{L}$ is a Hermitian preserving map, $\mathcal{P}$ is closed under taking adjoints. Now, to prove it is closed under multiplication, let $X,Y\in\mathbb{M}_{d}$ and $a,b\in\mathbb{R}$ be such that $\mathcal{L}(X)=iaX$ and $\mathcal{L}(Y)=ibY$. Then for all $t\geq 0$,
@@
This shows that $X,Y$ is in the peripheral space of $e^{t\mathcal{L}}$. Since the peripheral space of $e^{t\mathcal{L}}$ lies in the multiplicative domain of $e^{t\mathcal{L}}$, we get that
@@
By differentiating both sides at $t=0$, we obtain $\mathcal{L}(XY)=i(a+b)XY$. This completes the proof. ∎
The algebra $\mathcal{P}$ is known as the peripheral Poisson boundary of $\mathcal{L}$, or more appropriately that of the associated quantum dynamical semigroup. For more on this topic see [6]. The following is a one parameter semigroup version of Theorem 3.1 of [7].
@@
    $\mathbb{M}_{d}=\mathcal{P}\oplus\mathcal{N}$, where $\mathcal{N}=\{Y:\lim_{t\to\infty}\Phi_{t}(Y)=0\}$;
@@
    $\Phi_{t}=\alpha_{t}+\beta_{t}$, where $\alpha_{t}:=\Phi_{t}\circ\mathbf{E}_{\mathcal{P}}:\mathbb{M}_{d}\to\mathcal{P}\subseteq\mathbb{M}_{d}$, where $\mathbf{E}_{\mathcal{P}}$ is the unique trace-preserving conditional expectation onto $\mathcal{P}$ and $\beta_{t}:=\Phi_{t}-\alpha_{t}:\mathbb{M}_{d}\to\mathcal{N}\subseteq\mathbb{M}_{d}$ satisfies $\lim_{t\to\infty}\beta_{t}=0$.
@@
$(i)$ Using the Jordan decomposition of the generator $\mathcal{L}$ of $\{\Phi_{t}\}_{tgeq0}$, we write $\mathbb{M}_{d}$ as a vector space direct sum:
@@
where $\sigma(\mathcal{L})=\{\lambda_{1},\lambda_{2},\cdots,\lambda_{n}\}$ is the set of eigenvalues of $\mathcal{L}$ and $\nu(\lambda_{j})$ is the algebraic multiplicity of the eigenvalue $\lambda_{j}$. From Lemma 4.4, if the real part of $\lambda_{j}=0$, all the Jordan blocks corresponding to the eigenvalue $\lambda_{j}$ are of dimension one and hence the algebraic multiplicity is the same as its geometric multiplicity. Therefore,
@@
where $\sigma_{pi}(\mathcal{L})$ be the set of all purely imaginary eigenvalues of $\mathcal{L}$. Suppose $\lambda_{k}\in\sigma(\mathcal{L})$ be such that the real part of $\lambda_{k}\neq 0$, then again by Lemma 4.4, we have $e^{tJ_{d_{k}}(\lambda_{k})}\to 0$ as $t\to\infty$. Thus, for $Y\in\textsf{ker}(\mathcal{L}-\lambda_{k}\operatorname{\textsf{id}})^{\nu(\lambda_{k})}$, we have $\Phi_{t}(Y)\to 0$ as $t\to\infty$. By letting $\mathcal{N}:=\bigoplus_{\lambda\in\sigma(\mathcal{L})\setminus\sigma_{pi}(\mathcal{L})}\textsf{ker}(\mathcal{L}-\lambda\operatorname{\textsf{id}})^{\nu(\lambda)}$, we get the required decomposition. Note that as $\Phi_{t}$ is contractive in Hilbert-Schmidt norm and $\mathcal{P}$ is the span of peripheral eigenvectors, it is reducing for $\Phi_{t}$. Consequently, $\mathbb{M}_{d}=\mathcal{P}\oplus\mathcal{N}$, is an orthogonal direct sum with respect to the Hilbert-Schmidt inner product.
@@
$(iii)$ For every $t\geq 0$, let $\alpha_{t}:=\Phi_{t}\circ\mathbf{E}_{\mathcal{P}}$ and $\beta_{t}:=\Phi_{t}-\alpha_{t}$. Given $Z\in\mathbb{M}_{d}$, decompose it as $Z=X+Y$ with $X\in\mathcal{P}$ and $Y\in\mathcal{N}$. Then for all $t\geq 0$,
$\beta_{t}(Z)=\Phi_{t}(Z)-\alpha_{t}(Z)=\Phi_{t}(X)+\Phi_{t}(Y)-\Phi_{t}\circ\mathbf{E}_{\mathcal{P}}(Z)=\Phi_{t}(Y).$
Since $\Phi_{t}(Y)\to 0$ as $t\to\infty$, we have $\beta_{t}(\mathbb{M}_{d})\subseteq\mathcal{N}$ for all $t\geq 0$ and $\left\lVert\Phi_{t}-\alpha_{t}\right\rVert=\left\lVert\beta_{t}\right\rVert\to 0$ as $t\to\infty$. ∎
Recall that decoherence-free *-subalgebra of the dynamical semigroup $\{\Phi_{t}\}_{t\geq 0}$ of unital channels is the largest *-algebra $\mathcal{D}$ such that for every $X\in\mathcal{D}$ we have
$\Phi_{t}(X^{*}X)=\Phi_{t}(X^{*})\Phi_{t}(X)\qquad\text{and}\qquad\Phi_{t}(XX^{*})=\Phi_{t}(X)\Phi_{t}(X^{*}),\forall t\geq 0.$
Therefore, $\mathcal{D}=\bigcap_{t\geq 0}\mathscr{M}_{\Phi_{t}}$, where $\mathscr{M}_{\Phi}$ denotes the multiplicative domain of $\Phi$. Furthermore, if the generator $\mathcal{L}$ is given by 4.1, then from [13, 33], $\mathcal{D}$ is equal to the commutator of the set
$\{L_{j},L_{j}^{*},[H,L_{j}],[H,L_{j}^{*}],[H,[H,L_{j}]],[H,[H,L_{j}^{*}]],\cdots|1\leq j\leq n\}.$
The results in [13, 33] are applicable for quantum dynamical semigroups on $\mathscr{B}(\mathcal{H})$, where $\mathcal{H}$ is a separable Hilbert space, possibly infinite dimensional. We wish to show that in our context of unital quantum channels the decoherence-free *-subalgebra is nothing but the peripheral Poisson boundary and can be described in simpler ways.
Lemma 4.7.